I need to see how many solutions exist in this linear equation. Then I need to graph the two lines the linear equation makes. 0.2x + 0.4y = 1.7 and 8.3x - 6.3y = -4.3
Answer by stanbon(75887)

I need to see how many solutions exist in this linear equation. Then I need to graph the two lines the linear equation makes.
0.2x + 0.4y = 1.7
8.3x - 6.3y = -4.3
------------------------
Multiply thru both equations by 10 to get rid of the decimals:
2x + 4y = 17
83x - 63y = -43
-----------------------------
Multiply thru the 1st equation by 83; Multiply thru the 2nd equationby 2:
166x + 332y = 83*17
166x - 126y = -86
----------------------
Subtract the 2nd equation from the 1st to solve for "y":
458y = 1497
y = 3.268559..
----
Substitute into 2x + 4y = 17 to solve for "x":
x = 1.963495..
-----------------------

----------------------
Cheers,
Stan H.
y =